OpenStudy (aprilrm21):

PRECALC HELP PLEASE!!! Verify: csc^2xcot^2x + csc^2x = csc^4x

OpenStudy (aprilrm21):

@TheSmartOne

OpenStudy (welshfella):

i think we can use the identity csx^2 x = 1 + cot^2 x

OpenStudy (welshfella):

from this cot^2 x = csc^2 x - 1 plug this into the left side of the identity and simplify.

OpenStudy (aprilrm21):

So change all the csc to 1+cot^2x ?

OpenStudy (welshfella):

No change the cot^2 x to csc^2 x - 1

OpenStudy (aprilrm21):

So then would the exponents multiply in in the first part?

OpenStudy (welshfella):

so csc^2 x ( csc^2 x - 1) + csc^2 x Simplify this and see what you get.

OpenStudy (welshfella):

when you multiply you add exponents

OpenStudy (welshfella):

its now easy to solve.

OpenStudy (aprilrm21):

So right now I have csc^4x - 1 + csc^2x

OpenStudy (aprilrm21):

Am I on the right track?

OpenStudy (welshfella):

you made 1 mistake csc^2 x is distributed over every term in the parentheses.

OpenStudy (aprilrm21):

OH! So the last 2 terms cancle!

OpenStudy (welshfella):

you did csc^2 x + csc^2 x = csc^4 x but also its csc^2 x * -1

OpenStudy (welshfella):

yes they cancel leaving just csc^4 x

OpenStudy (aprilrm21):

Thank you so much!

OpenStudy (welshfella):

yw